OpenStudy (anonymous):

the square root of 512ksquared

OpenStudy (antonio_xx2):

16k * SQRT(2).

OpenStudy (anonymous):

\[\sqrt{512k^2}=\sqrt{512}\sqrt{k^2} = \sqrt{256\times2}\sqrt{k^2}=16\sqrt{2}\times k=16k \sqrt{2}\]